# Strong primeness for equivalence relations arising from Zariski dense subgroups

### v1

@@ Abstract.
CH is supported by Institut Universitaire de France
@@ Introduction
We say that a type ${\rm II_{1}}$ ergodic equivalence relation $\mathscr{R}$ is prime if for any ergodic equivalence relations $\mathscr{T}_{1},\mathscr{T}_{2}$ for which $\mathscr{R}\cong\mathscr{T}_{1}\times\mathscr{T}_{2}$, then there exists $i\in\{1,2\}$ such that $\mathscr{T}_{i}$ has finite orbits almost everywhere. Zimmer’s seminal work [Zi81] shows that for any lattice $\Gamma<G$ in a noncompact simple connected real Lie group with finite center and any essentially free ergodic pmp action $\Gamma\curvearrowright(X,\nu)$, the orbit equivalence relation $\mathscr{R}(\Gamma\curvearrowright X)$ is prime.
@@
In order to motivate and state our main results, we introduce the following terminology using the framework of Popa’s intertwining theory for equivalence relations (see Section 2 for further details). We say that a type ${\rm II_{1}}$ ergodic equivalence relation $\mathscr{R}$ is strongly prime if for any ergodic equivalence relations $\mathscr{S}$, $\mathscr{T}_{1}$, $\mathscr{T}_{2}$ for which ×$\mathscr{R}\times\mathscr{S}\cong\mathscr{T}_{1}\times\mathscr{T}_{2}$, then there exists $i\in\{1,2\}$ such that as subequivalence relations, $\mathscr{T}_{i}$ embeds into $\mathscr{S}$ inside ×$\mathscr{R}\times\mathscr{S}$, that we write $\mathscr{T}_{i}\preceq_{\mathscr{R}\times\mathscr{S}}\mathscr{S}$ (see Definition 2.4 for a precise statement). Note that this implies in particular that there exists an ergodic equivalence relation $\mathscr{U}$ such that $\mathscr{T}_{i}\times\mathscr{U}$ and $\mathscr{S}$ are stably isomorphic (see [Sp21]). We point out that the notion of strong primeness for type ${\rm II_{1}}$ factors was studied by Isono in [Is16].
@@
The ergodic theory of equivalence relations is intimately connected to the theory of von Neumann algebras. By a well-known construction due to Feldman–Moore [FM75], to any ergodic equivalence relation $\mathscr{R}$ on a standard probability space $(X,\nu)$, one can associate a von Neumann factor $\operatorname{L}(\mathscr{R})$ so that $\operatorname{L}^{\infty}(X,\nu)\subset\operatorname{L}(\mathscr{R})$ sits as a Cartan subalgebra. Recall that a type ${\rm II_{1}}$ factor $M$ is prime if for any factors $M_{1},M_{2}$ for which $M\cong M_{1}\mathbin{\overline{\otimes}}M_{2}$, then there exists $i\in\{1,2\}$ such that $M_{i}$ is finite dimensional. For any ergodic equivalence relation $\mathscr{R}$, if $\operatorname{L}(\mathscr{R})$ is prime, then $\mathscr{R}$ is prime, but the converse need not hold in general (see Proposition 3.7 below). In relation with Zimmer’s primeness result [Zi81], in the case of simple connected real Lie groups with finite center and real rank one, any discrete subgroup $\Gamma<G$ is biexact in the sense of [BO08]. Then by Ozawa’s relative solidity result [Oz04], it follows that for any essentially free ergodic pmp action $\Gamma\curvearrowright(X,\nu)$ of a countable discrete nonamenable biexact group, the group measure space factor $\operatorname{L}(\Gamma\curvearrowright X)$ is prime. For other iconic primeness results in von Neumann algebra theory, we refer to [Ge96, Oz03, Pe06].
@@
    If $\mathscr{R}=\mathscr{R}_{1}\times\cdots\times\mathscr{R}_{n}\cong\mathscr{T}_{1}\times\mathscr{T}_{2}$, where $\mathscr{T}_{1},\mathscr{T}_{2}$ are type ${\rm II_{1}}$ ergodic equivalence relations, then there exists a partition $\{1,\dots,n\}=T_{1}\sqcup T_{2}$ into nonempty subsets such that for every $j\in\{1,2\}$, we have $\mathscr{T}_{j}\preceq_{\mathscr{R}}\mathscr{R}_{T_{j}}$ as subequivalence relations, where $\mathscr{R}_{T_{j}}=\prod_{i\neq j}\mathscr{R}_{i}$.
@@ 2.1. Equivalence relations
We denote by $\operatorname{Aut}(X,\nu)$ the group of all pmp Borel automorphisms $\theta:(X,\nu)\to(X,\nu)$. We denote by $\operatorname{Aut}(\mathscr{R})$ the automorphism group of $\mathscr{R}$ that consists of all pmp Borel automorphisms $\theta\in\operatorname{Aut}(X,\nu)$ such that $(\theta(x),\theta(y))\in\mathscr{R}$ if and only if $(x,y)\in\mathscr{R}$ for $(x,y)\in X\times X$. Then we denote by $[\mathscr{R}]<\operatorname{Aut}(\mathscr{R})$ the full group of $\mathscr{R}$ that consists of all automorphisms $\theta\in\operatorname{Aut}(\mathscr{R})$ for which $(\theta(x),x)\in\mathscr{R}$ for $\nu$-almost every $x\in X$. The uniform metric
@@
Whenever $U\subset X$ is a nonnull measurable subset, we define the equivalence relation $\mathscr{R}|_{U}=\mathscr{R}\cap(U\times U)$ on the standard probability space $(U,\nu_{U})$ where $\nu_{U}=\frac{1}{\nu(U)}\nu|_{U}$. Assume that $\mathscr{R}$ is ergodic. For every $t>0$, we define the amplification $\mathscr{R}^{t}$ as follows. Denote by $c$ the counting measure on $\mathbb{N}$. Choose a measurable subset $X_{t}\subset X\times\mathbb{N}$ such that $(\nu\otimes c)(X_{t})=t$ and set $\nu_{t}=\frac{1}{(\nu\otimes c)(X_{t})}(\nu\otimes c)|_{X_{t}}$. Define the ergodic equivalence relation $\mathscr{R}^{t}$ on the standard probability space $(X_{t},\nu_{t})$ by declaring $((x,p),(y,q))\in\mathscr{R}^{t}$ if and only if $(x,y)\in\mathscr{R}$ for $((x,p),(y,q))\in X_{t}\times X_{t}$.
@@
Let $\mathscr{S}$ be another equivalence relation defined on a standard probability space $(Y,\eta)$. We say that $\mathscr{R}$ and $\mathscr{S}$ are isomorphic if there exists a pmp Borel isomorphism $\theta:(X,\nu)\to(Y,\eta)$ such that for $m$-almost every $(x,y)\in\mathscr{R}$ we have $(\theta(x),\theta(y))\in\mathscr{S}$. We say that $\mathscr{R}$ and $\mathscr{S}$ are stably isomorphic if there exist nonnull measurable subsets $U\subset X$ and $V\subset Y$ and a pmp Borel isomorphism $\theta:(U,\nu_{U})\to(V,\eta_{V})$ such that for $m$-almost every $(x,y)\in\mathscr{R}|_{U}$, we have $(\theta(x),\theta(y))\in\mathscr{S}|_{V}$.
@@ 2.2. Popa’s intertwining theory for equivalence relations
Let $\mathscr{R}$ be an ergodic equivalence relation. We say that $\mathscr{R}$ is strongly prime if for any ergodic equivalence relations $\mathscr{S}$, $\mathscr{T}_{1}$, $\mathscr{T}_{2}$ for which ×$\mathscr{R}\times\mathscr{S}\cong\mathscr{T}_{1}\times\mathscr{T}_{2}$, then there exists $i\in\{1,2\}$ such that we have $\mathscr{T}_{i}\preceq_{\mathscr{R}\times\mathscr{S}}\mathscr{S}$ as subequivalence relations.
@@ 2.5. Chabauty topology and Wijsman topology
Let $\mathscr{G}$ be a Polish group and choose a compatible complete metric $d:\mathscr{G}\times\mathscr{G}\to\mathbb{R}_{+}$. Denote by $\operatorname{CL}(\mathscr{G})$ the space of all nonempty closed subsets of $\mathscr{G}$. For all $x\in\mathscr{G}$, $A\in\operatorname{CL}(\mathscr{G})$, set $d(x,A)=\inf\left\{d(x,a)\mid a\in A\right\}$. Following [Be90], the Wijsman topology $\tau_{W(d)}$ on $\operatorname{CL}(\mathscr{G})$ is defined as the weakest topology on $\operatorname{CL}(\mathscr{G})$ that makes the maps $\operatorname{CL}(\mathscr{G})\to\mathbb{R}_{+}:A\mapsto d(x,A)$ continuous for all $x\in\mathscr{G}$. As explained in [Be90], the Borel structure induced by the Wijsman topology $\tau_{W(d)}$ coincides with the Effros-Borel structure on $\operatorname{CL}(\mathscr{G})$. By [Be90, Theorem 4.3], $(\operatorname{CL}(\mathscr{G}),\tau_{W(d)})$ is a Polish space. Consider the space $\operatorname{Sub}(\mathscr{G})$ of all closed subgroups of $\mathscr{G}$. Then $\operatorname{Sub}(\mathscr{G})\subset\operatorname{CL}(\mathscr{G})$ is closed with respect to $\tau_{W(d)}$ and so $(\operatorname{Sub}(\mathscr{G}),\tau_{W(d)})$ is a Polish space. Observe that the subset
@@ 3. Proofs of the main results
is isometric (hence injective) and surjective. Therefore, we may and will identify $\operatorname{L}^{0}(X\times Y,\nu\otimes\eta,\mathbb{C})$ and $\operatorname{L}^{0}(X,\nu,\operatorname{L}^{0}(Y,\eta,\mathbb{C}))$.
@@ 3.1. Zimmer’s method for product equivalence relations
Let $\mathscr{U}$ be an ergodic equivalence relation on $(U,\nu)$ and $\mathscr{V}$ an ergodic equivalence relation on $(V,\eta)$. Set $\mathscr{W}=\mathscr{U}\times\mathscr{V}$ and $(W,\zeta)=(U\times V,\nu\otimes\eta)$. Let $\alpha:\mathscr{W}\to G$ be a Borel $1$-cocycle. Following [Zi81], we define the set $\mathfrak{A}(\alpha)$ that consists of all pairs $(\varphi,\mathfrak{H})$, where $\varphi:W\to G$ is a Borel map, $\mathfrak{H}:U\to\operatorname{Sub}(G):u\mapsto H_{u}$ is a Borel map such that for $\nu$-almost every $u\in U$, there exists a $k$-subgroup $\mathbf{H}_{u}<\mathbf{G}$ such that $\mathbf{H}_{u}(k)=H_{u}$ and the Borel map $V\to G/H_{u}:v\mapsto\varphi(u,v)H_{u}$ is $(\mathscr{V},\alpha_{u})$-equivariant, where $\alpha_{u}:\mathscr{V}\to G:(v_{1},v_{2})\mapsto\alpha((u,v_{1}),(u,v_{2}))$. Observe that the trivial pair $(\varphi,\mathfrak{H})$ defined by $\varphi:W\to G:(u,v)\mapsto e$ and $\mathfrak{H}:U\to\operatorname{Sub}(G):u\mapsto G$ belongs to $\mathfrak{A}(\alpha)$. Thus, $\mathfrak{A}(\alpha)$ is not the empty set.
As explained in [Zi81], the set $\mathfrak{A}(\alpha)$ is stable under the following cutting and pasting procedures. Firstly, let $(\varphi,\mathfrak{H})\in\mathfrak{A}(\alpha)$ and $\lambda\in[\mathscr{U}]$. Define the pair $(\psi,\mathfrak{J})=\lambda_{\ast}(\varphi,\mathfrak{H})$ by $\psi:W\to G:(u,v)\mapsto\alpha((u,v),(\lambda u,v))\varphi(\lambda u,v)$ and $\mathfrak{J}:U\to\operatorname{Sub}(G):u\mapsto H_{\lambda u}$. Then a straightforward computation shows that $(\psi,\mathfrak{J})\in\mathfrak{A}(\alpha)$. Secondly, let $(\varphi^{i},\mathfrak{H}^{i})\in\mathfrak{A}(\alpha)$ for every $i\in\{1,2\}$. Let $U=U_{1}\sqcup U_{2}$ be a measurable partition. Define the pair $(\varphi,\mathfrak{H})$ by $\varphi|_{U_{i}\times V}=\varphi^{i}|_{U_{i}\times V}$ and $\mathfrak{H}|_{U_{i}\times V}=\mathfrak{H}^{i}|_{U_{i}\times V}$ for every $i\in\{1,2\}$. Then it is plain to see that $(\varphi,\mathfrak{H})\in\mathfrak{A}(\alpha)$. We record the following useful result.
@@
There exists a pair $(\varphi,\mathfrak{H})\in\mathfrak{A}(\alpha)$ such that the $\nu$-almost everywhere defined map $U\mapsto\mathbb{N}:u\mapsto\dim(\mathbf{H}_{u})$ is $\nu$-almost everywhere constant and for any other pair $(\psi,\mathfrak{J})\in\mathfrak{A}(\alpha)$, we have $\dim(\mathbf{J}_{u})\geq\dim(\mathbf{H}_{u})$ for $\nu$-almost every $u\in U$.
@@
Keep the same notation as above. The following assertions hold:
@@
$(\rm i)$ After discarding a $\nu$-null Borel subset, we may assume that the Borel maps $\mathfrak{H}:U\to\operatorname{Sub}(G):u\mapsto H_{u}$ and $\mathfrak{J}:U\to\operatorname{Sub}(G):u\mapsto J_{u}$ satisfy that for every $u\in U$, there exist a $k$-subgroup $\mathbf{H}_{u}<\mathbf{G}$ such that $\mathbf{H}_{u}(k)=H_{u}$ and a $k$-subgroup $\mathbf{J}_{u}<\mathbf{G}$ such that $\mathbf{J}_{u}(k)=J_{u}$, and the Borel maps $V\to G/H_{u}:v\mapsto\varphi(u,v)H_{u}$ and $V\to G/J_{u}:v\mapsto\psi(u,v)J_{u}$ are $(\mathscr{V},\alpha_{u})$-equivariant. Denote by $\mathscr{G}=\operatorname{L}^{0}(V,\eta,G)$ the space of all $\eta$-equivalence classes of measurable functions $f:V\to G$. Endowed with the topology of convergence in measure, $\mathscr{G}$ is a Polish group. For every $u\in U$, consider the $k$-$\mathbf{G}$-variety $\mathbf{M}_{u}=\mathbf{G}/\mathbf{H}_{u}\times\mathbf{G}/\mathbf{J}_{u}$, where $\mathbf{G}\curvearrowright\mathbf{M}_{u}$ acts diagonally. Set $M_{u}=\mathbf{M}_{u}(k)$ and regard $N_{u}=G/H_{u}\times G/J_{u}\subset M_{u}$ as a closed subset. Define the Borel map
@@
is a homeomorphism. By considering a Borel section $G/L\to G$ of the projection map $G\to G/L$, we may then find $F\in\mathscr{G}$ such that $\theta_{u}(v)=(F(v)H_{u},F(v)bJ_{u})$ for $\eta$-almost every $v\in V$.
For every $u\in U$, define $\varphi_{u}:V\to G:v\mapsto\varphi(u,v)$, $\psi_{u}:V\to G:v\mapsto\psi(u,v)$ and regard $\varphi_{u}\in\mathscr{G}$, $\psi_{u}\in\mathscr{G}$. Then the maps $U\to\mathscr{G}:u\mapsto\varphi_{u}$ and $U\to\mathscr{G}:u\mapsto\psi_{u}$ are Borel. For every $u\in U$, set $\mathscr{H}_{u}=\operatorname{L}^{0}(V,\eta,H_{u})$ and $\mathscr{J}_{u}=\operatorname{L}^{0}(V,\eta,J_{u})$. By Proposition 2.9, the maps $U\to\operatorname{Sub}(\mathscr{G}):u\mapsto\mathscr{H}_{u}$ and $U\to\operatorname{Sub}(\mathscr{G}):u\mapsto\mathscr{J}_{u}$ are Borel. Denote by $\pi_{U}:U\times G\times\mathscr{G}\to U$ the projection map. Consider the subset

### v2

@@ Abstract.
CH is supported by Institut Universitaire de France and ERC Advanced Grant NET 101141693
@@ Introduction
We say that a type ${\rm II_{1}}$ ergodic equivalence relation $\mathscr{R}$ is prime when for any ergodic equivalence relations $\mathscr{T}_{1},\mathscr{T}_{2}$ for which $\mathscr{R}\cong\mathscr{T}_{1}\times\mathscr{T}_{2}$, there exists $i\in\{1,2\}$ such that $\mathscr{T}_{i}$ has finite orbits almost everywhere. Zimmer’s seminal work [Zi81] shows that for any lattice $\Gamma<G$ in a noncompact simple connected real Lie group with finite center and any essentially free ergodic pmp action $\Gamma\curvearrowright(X,\nu)$, the orbit equivalence relation $\mathscr{R}(\Gamma\curvearrowright X)$ is prime.
@@
In order to motivate and state our main results, we introduce the following terminology using the framework of Popa’s intertwining theory for equivalence relations (see Section 2 for further details). We say that a type ${\rm II_{1}}$ ergodic equivalence relation $\mathscr{R}$ is strongly prime when for any ergodic equivalence relations $\mathscr{S}$, $\mathscr{T}_{1}$, $\mathscr{T}_{2}$ for which ×$\mathscr{R}\times\mathscr{S}\cong\mathscr{T}_{1}\times\mathscr{T}_{2}$, there exists $i\in\{1,2\}$ such that as subequivalence relations, $\mathscr{T}_{i}$ embeds into $\mathscr{S}$ inside ×$\mathscr{R}\times\mathscr{S}$, that we write $\mathscr{T}_{i}\preceq_{\mathscr{R}\times\mathscr{S}}\mathscr{S}$ (see Definition 2.4 for a precise statement). Note that this implies in particular that there exists an ergodic equivalence relation $\mathscr{U}$ such that $\mathscr{T}_{i}\times\mathscr{U}$ and $\mathscr{S}$ are stably isomorphic (see [Sp21]). We point out that the notion of strong primeness for type ${\rm II_{1}}$ factors was studied by Isono in [Is16].
@@
The ergodic theory of equivalence relations is intimately connected to the theory of von Neumann algebras. By a well-known construction due to Feldman–Moore [FM75], to any ergodic equivalence relation $\mathscr{R}$ on a standard probability space $(X,\nu)$, one can associate a von Neumann factor $\operatorname{L}(\mathscr{R})$ so that $\operatorname{L}^{\infty}(X,\nu)\subset\operatorname{L}(\mathscr{R})$ sits as a Cartan subalgebra. Recall that a type ${\rm II_{1}}$ factor $M$ is prime when for any factors $M_{1},M_{2}$ for which $M\cong M_{1}\mathbin{\overline{\otimes}}M_{2}$, there exists $i\in\{1,2\}$ such that $M_{i}$ is finite dimensional. For any ergodic equivalence relation $\mathscr{R}$, if $\operatorname{L}(\mathscr{R})$ is prime, then $\mathscr{R}$ is prime, but the converse need not hold in general (see Proposition 3.7 below). In relation with Zimmer’s primeness result [Zi81], in the case of simple connected real Lie groups with finite center and real rank one, any discrete subgroup $\Gamma<G$ is biexact in the sense of [BO08]. Then by Ozawa’s relative solidity result [Oz04], it follows that for any essentially free ergodic pmp action $\Gamma\curvearrowright(X,\nu)$ of a countable discrete nonamenable biexact group, the group measure space factor $\operatorname{L}(\Gamma\curvearrowright X)$ is prime. For other iconic primeness results in von Neumann algebra theory, we refer to [Ge96, Oz03, Pe06].
@@
    If $\mathscr{R}=\mathscr{R}_{1}\times\cdots\times\mathscr{R}_{n}\cong\mathscr{T}_{1}\times\mathscr{T}_{2}$, where $\mathscr{T}_{1},\mathscr{T}_{2}$ are type ${\rm II_{1}}$ ergodic equivalence relations, then there exists a partition $\{1,\dots,n\}=T_{1}\sqcup T_{2}$ into nonempty subsets such that for every $j\in\{1,2\}$, we have $\mathscr{T}_{j}\preceq_{\mathscr{R}}\mathscr{R}_{T_{j}}$ as subequivalence relations, where $\mathscr{R}_{T_{j}}=\prod_{i\in T_{j}}\mathscr{R}_{i}$.
@@
The statement of Corollary C$(\rm ii)$ is sharper than the one in Corollary B$(\rm ii)$ in the sense that we may choose $t_{1},\dots,t_{n}>0$ such that $\mathscr{T}_{i}^{t_{i}}\cong\mathscr{R}_{i}$ for every $1\leq i\leq n$ in such a way that $t_{1}\cdots t_{n}=1$.
@@ 2.1. Equivalence relations
We denote by $\operatorname{Aut}(X,\nu)$ the group of all pmp Borel automorphisms $\theta:(X,\nu)\to(X,\nu)$. We denote by $\operatorname{Aut}(\mathscr{R})$ the automorphism group of $\mathscr{R}$ that consists of all pmp Borel automorphisms $\theta\in\operatorname{Aut}(X,\nu)$ such that $(\theta(x),\theta(y))\in\mathscr{R}$ and $(\theta^{-1}(x),\theta^{-1}(y))\in\mathscr{R}$ for $m$-almost every $(x,y)\in\mathscr{R}$. Then we denote by $[\mathscr{R}]<\operatorname{Aut}(\mathscr{R})$ the full group of $\mathscr{R}$ that consists of all automorphisms $\theta\in\operatorname{Aut}(\mathscr{R})$ for which $(\theta(x),x)\in\mathscr{R}$ for $\nu$-almost every $x\in X$. The uniform metric
@@
Whenever $U\subset X$ is a nonnull measurable subset, we define the equivalence relation $\mathscr{R}|_{U}=\mathscr{R}\cap(U\times U)$ on the standard probability space $(U,\nu_{U})$ where $\nu_{U}=\frac{1}{\nu(U)}\nu|_{U}$. Assume that $\mathscr{R}$ is ergodic and that $(X,\nu)$ is diffuse. For every $t>0$, we define the amplification $\mathscr{R}^{t}$ as follows. Denote by $c$ the counting measure on $\mathbb{N}$. Choose a measurable subset $X_{t}\subset X\times\mathbb{N}$ such that $(\nu\otimes c)(X_{t})=t$ and set $\nu_{t}=\frac{1}{(\nu\otimes c)(X_{t})}(\nu\otimes c)|_{X_{t}}$. Define the ergodic equivalence relation $\mathscr{R}^{t}$ on the standard probability space $(X_{t},\nu_{t})$ by declaring $((x,p),(y,q))\in\mathscr{R}^{t}$ if and only if $(x,y)\in\mathscr{R}$ for $((x,p),(y,q))\in X_{t}\times X_{t}$.
@@
Let $\mathscr{S}$ be another equivalence relation defined on a standard probability space $(Y,\eta)$. We say that $\mathscr{R}$ and $\mathscr{S}$ are isomorphic if there exists a pmp Borel isomorphism $\theta:(X,\nu)\to(Y,\eta)$ such that for almost every $(x,y)\in\mathscr{R}$, we have $(\theta(x),\theta(y))\in\mathscr{S}$ and for almost every $(x,y)\in\mathscr{S}$, we have $(\theta^{-1}(x),\theta^{-1}(y))\in\mathscr{R}$. Assuming moreover that $\mathscr{R}$ and $\mathscr{S}$ are both ergodic, we say that $\mathscr{R}$ and $\mathscr{S}$ are stably isomorphic if there exist nonnull measurable subsets $U\subset X$ and $V\subset Y$ and a pmp Borel isomorphism $\theta:(U,\nu_{U})\to(V,\eta_{V})$ such that for almost every $(x,y)\in\mathscr{R}|_{U}$, we have $(\theta(x),\theta(y))\in\mathscr{S}|_{V}$ and for almost every $(x,y)\in\mathscr{S}|_{V}$, we have $(\theta^{-1}(x),\theta^{-1}(y))\in\mathscr{R}|_{U}$.
@@ 2.2. Popa’s intertwining theory for equivalence relations
Let $\mathscr{R}$ be an ergodic equivalence relation. We say that $\mathscr{R}$ is strongly prime when for any ergodic equivalence relations $\mathscr{S}$, $\mathscr{T}_{1}$, $\mathscr{T}_{2}$ for which ×$\mathscr{R}\times\mathscr{S}\cong\mathscr{T}_{1}\times\mathscr{T}_{2}$, there exists $i\in\{1,2\}$ such that we have $\mathscr{T}_{i}\preceq_{\mathscr{R}\times\mathscr{S}}\mathscr{S}$ as subequivalence relations.
@@ 2.5. Chabauty topology and Wijsman topology
Let $\mathscr{G}$ be a Polish group and choose a compatible complete metric $d:\mathscr{G}\times\mathscr{G}\to\mathbb{R}_{+}$. Denote by $\operatorname{CL}(\mathscr{G})$ the space of all nonempty closed subsets of $\mathscr{G}$. For all $x\in\mathscr{G}$, $A\in\operatorname{CL}(\mathscr{G})$, set $d(x,A)=\inf\left\{d(x,a)\mid a\in A\right\}$. Following [Be90], the Wijsman topology $\tau_{W(d)}$ on $\operatorname{CL}(\mathscr{G})$ is defined as the weakest topology on $\operatorname{CL}(\mathscr{G})$ that makes the maps $\operatorname{CL}(\mathscr{G})\to\mathbb{R}_{+}:A\mapsto d(x,A)$ continuous for all $x\in\mathscr{G}$. As explained in [Be90], the Borel structure induced by the Wijsman topology $\tau_{W(d)}$ coincides with the Effros–Borel structure on $\operatorname{CL}(\mathscr{G})$. By [Be90, Theorem 4.3], $(\operatorname{CL}(\mathscr{G}),\tau_{W(d)})$ is a Polish space. Consider the space $\operatorname{Sub}(\mathscr{G})$ of all closed subgroups of $\mathscr{G}$. Then $\operatorname{Sub}(\mathscr{G})\subset\operatorname{CL}(\mathscr{G})$ is closed with respect to $\tau_{W(d)}$ and so $(\operatorname{Sub}(\mathscr{G}),\tau_{W(d)})$ is a Polish space. Observe that the subset
@@ 3. Proofs of the main results
is isometric (hence injective) and surjective. Therefore, we may and will identify $\operatorname{L}^{0}(X\times Y,\nu\otimes\eta,\mathbb{C})$ and $\operatorname{L}^{0}(X,\nu,\operatorname{L}^{0}(Y,\eta,\mathbb{C}))$ (see e.g. [Mo75, Theorem 1]).
@@ 3.1. Zimmer’s method for product equivalence relations
Let $\mathscr{U}$ be an ergodic equivalence relation on $(U,\nu)$ and $\mathscr{V}$ an ergodic equivalence relation on $(V,\eta)$. Set $\mathscr{W}=\mathscr{U}\times\mathscr{V}$ and $(W,\zeta)=(U\times V,\nu\otimes\eta)$. Let $\alpha:\mathscr{W}\to G$ be a Borel $1$-cocycle. Following [Zi81], we define the set $\mathfrak{A}(\alpha)$ that consists of all pairs $(\varphi,\mathfrak{H})$, where $\varphi:W\to G$ is a Borel map, $\mathfrak{H}:U\to\operatorname{Sub}(G):u\mapsto H_{u}$ is a Borel map such that for $\nu$-almost every $u\in U$, there exists a $k$-subgroup $\mathbf{H}_{u}<\mathbf{G}$ such that $\mathbf{H}_{u}(k)=H_{u}$ and the Borel map $V\to G/H_{u}:v\mapsto\varphi(u,v)H_{u}$ is $(\mathscr{V},\alpha_{u})$-equivariant, where $\alpha_{u}:\mathscr{V}\to G:(v_{1},v_{2})\mapsto\alpha((u,v_{1}),(u,v_{2}))$. Observe that the trivial pair $(\varphi,\mathfrak{H})$ defined by $\varphi:W\to G:(u,v)\mapsto e$ and $\mathfrak{H}:U\to\operatorname{Sub}(G):u\mapsto G$ belongs to $\mathfrak{A}(\alpha)$. Thus, $\mathfrak{A}(\alpha)$ is not the empty set. We note that the set $\mathfrak{A}(\alpha)$ only depends on the restriction of $\alpha$ to $\operatorname{id}_{U}\times\mathscr{V}$.
As explained in [Zi81], the set $\mathfrak{A}(\alpha)$ is stable under the following cutting and pasting procedures. To do this, we will exploit the fact that the Borel $1$-cocycle $\alpha$ is defined on $\mathscr{U}\times\mathscr{V}=\mathscr{W}$. Firstly, let $(\varphi,\mathfrak{H})\in\mathfrak{A}(\alpha)$ and $\lambda\in[\mathscr{U}]$. Define the pair $(\psi,\mathfrak{J})=\lambda_{\ast}(\varphi,\mathfrak{H})$ by $\psi:W\to G:(u,v)\mapsto\alpha((u,v),(\lambda u,v))\varphi(\lambda u,v)$ and $\mathfrak{J}:U\to\operatorname{Sub}(G):u\mapsto H_{\lambda u}$. Then a straightforward computation shows that $(\psi,\mathfrak{J})\in\mathfrak{A}(\alpha)$. Secondly, let $(\varphi^{i},\mathfrak{H}^{i})\in\mathfrak{A}(\alpha)$ for every $i\in\{1,2\}$. Let $U=U_{1}\sqcup U_{2}$ be a measurable partition. Define the pair $(\varphi,\mathfrak{H})$ by $\varphi|_{U_{i}\times V}=\varphi^{i}|_{U_{i}\times V}$ and $\mathfrak{H}|_{U_{i}\times V}=\mathfrak{H}^{i}|_{U_{i}\times V}$ for every $i\in\{1,2\}$. Then it is plain to see that $(\varphi,\mathfrak{H})\in\mathfrak{A}(\alpha)$. We record the following useful result.
@@
Assume that the probability space $(U,\nu)$ is nonatomic. There exists a pair $(\varphi,\mathfrak{H})\in\mathfrak{A}(\alpha)$ such that the $\nu$-almost everywhere defined map $U\mapsto\mathbb{N}:u\mapsto\dim(\mathbf{H}_{u})$ is $\nu$-almost everywhere constant and for any other pair $(\psi,\mathfrak{J})\in\mathfrak{A}(\alpha)$, we have $\dim(\mathbf{J}_{u})\geq\dim(\mathbf{H}_{u})$ for $\nu$-almost every $u\in U$.
@@
Assume that the probability space $(U,\nu)$ is nonatomic. Keep the same notation as above. The following assertions hold:
@@
$(\rm i)$ After discarding a $\nu$-null Borel subset, we may assume that the Borel maps $\mathfrak{H}:U\to\operatorname{Sub}(G):u\mapsto H_{u}$ and $\mathfrak{J}:U\to\operatorname{Sub}(G):u\mapsto J_{u}$ satisfy that for every $u\in U$, there exist a $k$-subgroup $\mathbf{H}_{u}<\mathbf{G}$ such that $\mathbf{H}_{u}(k)=H_{u}$ and a $k$-subgroup $\mathbf{J}_{u}<\mathbf{G}$ such that $\mathbf{J}_{u}(k)=J_{u}$, and the Borel maps $V\to G/H_{u}:v\mapsto\varphi(u,v)H_{u}$ and $V\to G/J_{u}:v\mapsto\psi(u,v)J_{u}$ are $(\mathscr{V},\alpha_{u})$-equivariant. Denote by $\mathscr{G}=\operatorname{L}^{0}(V,\eta,G)$ the space of all $\eta$-equivalence classes of measurable functions $f:V\to G$. Endowed with the topology of convergence in measure, $\mathscr{G}$ is a Polish group.
In this paragraph, we fix $u\in U$. Consider the $k$-$\mathbf{G}$-variety $\mathbf{M}_{u}=\mathbf{G}/\mathbf{H}_{u}\times\mathbf{G}/\mathbf{J}_{u}$, where $\mathbf{G}\curvearrowright\mathbf{M}_{u}$ acts diagonally. Set $M_{u}=\mathbf{M}_{u}(k)$ and regard $N_{u}=G/H_{u}\times G/J_{u}\subset M_{u}$ as a closed subset. Define the Borel map
@@
is a homeomorphism. By considering a Borel section $G/L\to G$ of the projection map $G\to G/L$, we may then find $F\in\mathscr{G}$ such that $\theta_{u}(v)=(F(v)H_{u},F(v)bJ_{u})$ for $\eta$-almost every $v\in V$. Define $\varphi_{u}:V\to G:v\mapsto\varphi(u,v)$, $\psi_{u}:V\to G:v\mapsto\psi(u,v)$ and regard $\varphi_{u}\in\mathscr{G}$, $\psi_{u}\in\mathscr{G}$.
Then the maps $U\to\mathscr{G}:u\mapsto\varphi_{u}$ and $U\to\mathscr{G}:u\mapsto\psi_{u}$ are Borel. For every $u\in U$, set $\mathscr{H}_{u}=\operatorname{L}^{0}(V,\eta,H_{u})$ and $\mathscr{J}_{u}=\operatorname{L}^{0}(V,\eta,J_{u})$. By Proposition 2.9, the maps $U\to\operatorname{Sub}(\mathscr{G}):u\mapsto\mathscr{H}_{u}$ and $U\to\operatorname{Sub}(\mathscr{G}):u\mapsto\mathscr{J}_{u}$ are Borel. Denote by $\pi_{U}:U\times G\times\mathscr{G}\to U$ the projection map. Consider the subset